Determine in each case whether the given planes are parallel or perpendicular:

1.x+y+3z+10=0 and x+2y-z=1,

2.3x-2y+z-6=0 and 4x+2y-4z=0,

3.3x+y+z-1=0 and -x+2y+z+3=0,

4.x-3y+z+1=0 and 3x-4y+z-1=0.


1
Expert's answer
2021-06-27T18:40:34-0400

1.x+y+3z+10=0 and x+2y-z=1


"\vec n_1=\langle1,1,3\rangle, \vec n_2=\langle1,2,-1\rangle""\dfrac{1}{1}\not=\dfrac{2}{1}"

The vectors "\vec n_1" and "\vec n_2" are not collinear.



"\vec n_1\cdot\vec n _2=1(1)+1(2)+3(-1)=0"

The vectors "\vec n_1" and "\vec n_2" are orthogonal.

Hence, the given planes are perpendicular.


2.3x-2y+z-6=0 and 4x+2y-4z=0


"\vec n_1=\langle3,-2,1\rangle, \vec n_2=\langle4,2,-4\rangle""\dfrac{4}{3}\not=\dfrac{2}{-2}"

The vectors "\vec n_1" and "\vec n_2" are not collinear.



"\vec n_1\cdot\vec n _2=3(4)+(-2)(2)+1(-4)=4\not=0"

The vectors "\vec n_1" and "\vec n_2" are not orthogonal.

Hence, the given planes are neither parallel nor perpendicular.


3.3x+y+z-1=0 and -x+2y+z+3=0


"\vec n_1=\langle3,1,1\rangle, \vec n_2=\langle-1,2,1\rangle""\dfrac{-1}{3}\not=\dfrac{2}{1}"

The vectors "\vec n_1" and "\vec n_2" are not collinear.



"\vec n_1\cdot\vec n _2=3(-1)+1(2)+1(1)=0"

The vectors "\vec n_1" and "\vec n_2" are orthogonal.

Hence, the given planes are perpendicular.


4.x-3y+z+1=0 and 3x-4y+z-1=0


"\vec n_1=\langle1,-3,1\rangle, \vec n_2=\langle3,-4,1\rangle""\dfrac{3}{1}\not=\dfrac{-4}{-3}"

The vectors "\vec n_1" and "\vec n_2" are not collinear.



"\vec n_1\cdot\vec n _2=1(3)+(-3)(-4)+1(1)=16\not=0"

The vectors "\vec n_1" and "\vec n_2" are not orthogonal.

Hence, the given planes are neither parallel nor perpendicular.